You have $30,000 to invest in a stock portfolio. Your choices are Stock X with an expected return of 14 percent and Stock Y with an expected return of 9 percent.

   

Required:
(a)

If your goal is to create a portfolio with an expected return of 12.4 percent, how much money will you invest in Stock X?

$74,400

$19,380

$21,216

$21,420

$20,400

  

(b)

If your goal is to create a portfolio with an expected return of 12.4 percent, how much money will you invest in Stock Y?

$9,600

$10,080

$9,216

$9,120

$9,984
Answer #1

Let investment in X=$x

Hence investment in Y=(30000-x)

Portfolio return=Respective return*Respective weight

12.4=(x/30000*14)+(30000-x)/30000*9

(12.4*30000)=14x+(30000-x)*9

372000=14x+270,000-9x

x=(372000-270,000)/(14-9)

=$20400=investment in X

Hence investment in Y=(30000-20400)=$9600.